# More than 200 Globular Clusters in the Milky Way, and still no super-Solar metallicity ones

### v1

@@ 4 Metallicity distribution
Moreover, comparing all distributions, we can see that the number density of the metal-poor part is in better agreement, whereas the metal-rich one is shifted to higher values for the new and bulge GCs samples than the other two. We can expect this effect because the metal-rich component is concentrated more towards the inner regions of the Galaxy, where the majority of our new targets are located, while the metal-poor one becomes more significant at larger Galactocentric distances. Indeed, as found by Tonini (2013) model (their Figure 2), the metal-rich distribution was formed locally in the main progenitor at $z\sim 2$, while the metal-poor component was formed in satellites at an epoch $z\sim 3-4$, and then merged with the main progenitor. Therefore, connecting to the discussion made in Section 3, and comparing our distribution with the Tonini’s model, the metallicity bimodaly is a direct prediction of a hierarchical clustering scenario. In consequence, we argue that the metal-poor component is mainly the contribution due to the tidal disruption of dwarf-like objects whereas the metal-rich population is related to the formation of the bulge and/or disk (see e.g. Mouhcine 2006). This may imply that many of our targets may be formed in-situ, and they might be classified as Main Progenitors following the Massari et al. (2019) nomenclature.
@@
Moreover, the height of the MD’s peaks can give us clues about the star formation histories. Figures 5 and 6 display that metal-poor and metal-rich peaks are similar, especially if we include in the total compilation the new GC sample. Therefore, based on the Tonini’s model, this trend suggests that the MW experienced a balanced star formation ($M_{SF}/M_{0}=0.2$) and merger history ($M_{1}/M_{0}=0.3$) [^6] Where $M_{0}$ is the final stellar mass, $M_{SF}/M_{0}$ star formation history parameter, and $M_{1}/M_{0}$ is assembly histories parameter.. Moreover, we know that multiple components, with different ages and metallicity, co-exist in the Bulge, suggesting that star formation does not stop at $z\sim 2$, but the final stellar mass is formed at later (or $z<2$) time, hinting that younger (e.g. age $\sim 10$ Gyr) GCs may have formed later in the MW bulge, due to an increasing in gas and so a burst in star formation due to merging events. Once the age of our targets has been confirmed, our young candidates might belong to this younger populations, and this can help to confirm if the formation of GCs can occur later in the MW bulge, indicating a strong presence of gas at that stage; otherwise this could be linked to accretion events, which enriched our Galaxy and formed a new generation of star clusters. However, this is a speculation, until this statement will be confirmed observationally and via simulations and models.
@@
Finally, we want to highlight the fact that spectroscopic analysis should be performed, in order to recover orbital parameters and chemical abundances, and thus constraining the real nature of these systems. Also, probably many other GCs are missing from our compilation.
@@
Figure 6: Similar to Figure 5 (right panel). We split the Harris (1996) (2010 edition) GCs catalogue into inner (bulge) GCs located at $R_{G}<3$ kpc (red line) and outer (halo) GCs at $R_{G}>3$ kpc (blue line). The new GCs sample is drawn in black line. We use this comparison to better understand the separation between bulge and halo clusters and to show that new GCs sample follow a similar trend as the inner GCs.
@@ 5 Comparisons with field stars
Another important comparison that we carried out is between GCs and FS MDs (see Figure 5). In order to do this, we took into consideration five Bulge FS public samples by Zoccali et al. (2008); Bensby et al. (2017); Hasselquist et al. (2020); Abdurro’uf et al. (2022). For the Abdurro’uf et al. (2022) catalogue, we applied the typical APOGEE selection cuts: apo_ps41 $<3.5$, ASPCAPFLAG $=0$, STARFLAG $=0$, SNR $>70$, LOGG $<3.6$, TEFF $<5500$, VSCATTER $<1$, and ruwe $<1.4$. Also, we included for comparison Galactic disk FS, using Casali et al. (2020) catalogue.
@@
Even though they come from widely different surveys (this was done intentionally since we need high statistics), these are all in good agreement, exhibiting metallicity peaks between -0.5 and 0.0, with a long tail towards the metal-poor regime. As well, we can see that both bulge and disk FS catalogues seems to follow the same behaviour. From Figure 5 we can see that $\sim 30-40\%$ of stars in all those catalogues show metallicities higher than Solar value, while a very low percentage of stars have [Fe/H] $<-1.0$. Even though we observe numerous very metal-rich stars, today we do not know any GC in the super-Solar metallicity regime. So, where are super metal-rich Bulge GCs? Did they ever exist? Will they exist? Can’t they even form? Is there a physical process that preferentially destroyed them in the past?
@@
We need to understand if there is a connection between GCs and isolated FS. We could resort to various possible explanations, such as a different star formation scenarios that may have occurred for these two populations as well as the formation of the bulge and disk during the early phases of the Universe (Costantin et al. 2021); or evolutionary effects, such as cluster disruption, mass loss, interactions, merging events (Minniti et al. 2017, 2019a,b). Likely, all of these have affected the ratio of clusters to FS at a given metallicity range and location within the Galaxy, and they may be responsible for these observed differences. Up to date, many studies were focused on the Galactic halo, finding a good agreement between observations and simulations. For instance, Reina-Campos et al. (2020), using E-MOSAIC simulation, found that GCs play a sub-dominant role in the build-up of the stellar halo. Additionally, they found that the disruption of GCs contributes between $0.3-14\%$ of the bulge mass, in agreement with observational estimates by Hughes et al. (2019b), which also predicted that stars from disrupted GCs in the Galactic bulge show the higher fraction around [Fe/H] $\approx-1$. A similar fraction was found by Schiavon et al. (2016), analysing FS in the MW bulge in a specific metallicity range of [Fe/H] $<-1$ and, by finding nitrogen-enriched stars, concluded that $14\%$ of the stellar mass of the bulge came from disrupted GCs. Therefore, the contribution of disrupted GCs to the stellar population of the bulge represent only a small percentage. Thus, we suspect that this factor alone cannot explain this inconsistency in metallicity between clusters and bulge FS, and probably the ratio of FS to GCs varies strongly in different environments. For instance, as found by Lamers et al. (2017) the difference between the GC and FS metallicity distributions is not specific to our Galaxy, but is a well-known phenomenon also in external galaxy, and it changes if outer or inner regions are taken into account. Reporting what they found: (i) in dwarf galaxies the population is in general metal-poor, but the GCs are usually more metal-poor than FS, (ii) in the M31 inner and outer halo, the cluster-to-star ratio is high at low metallicity and low at high metallicity, (iii) in both the inner and outer regions the relative contribution of the clusters at low metallicity is larger than that of the stars, (iv) in both the inner and the outer regions the relative contributions of the clusters at high metallicity is smaller than that of the stars, (v) finally, the ratios of Fornax dwarf spheroidal galaxy are more similar to those in the outer regions of the larger galaxies than to the inner regions.
5.1 Age–Metallicity distribution
@@
Another distribution that we want to take into consideration in order to understand the main differences between GCs and FS is the Age–Metallicity distribution (Figure 7). We used the catalogues by Bensby et al. (2017) (red points) and Casali et al. (2020) (cyan points), while we considered only the new GCs (orange points) by the present compilation, since in general these are low-luminosity and low-mass clusters, so we foresee that they have lost a fraction of stars during their life (see Section 3). As expected, we can see that there is an anti-correlation between age and metallicity, since the age decreases with metallicity. In particular, we can note that the younger and metal-rich population is represented by FS, while GCs occupy the right-bottom part of the diagram, with lower metallicities and older ages. This is also highlighted by the histograms and cumulative distributions displayed by Figure 7. GCs are more segregated at age $>10$ Gyr and their [Fe/H] spans over all values, especially at [Fe/H] $<-0.5$. In our compilation, the only two clusters with age $<10$ are VVVCL002 (age $>6.5$ Gyr) and Patchick 126 (age $>8$ Gyr). For the first, Minniti et al. (2023) demonstrated that VVVCL0002 is an old GC, since its high $\alpha$-enhancement ($[\alpha/Fe]>+0.4$) is connected to rapid chemical evolution dominated by core-collapse supernovae rather than type Ia supernovae, formed together with other clusters and field stars present today in the Galactic bulge, rather than a younger open cluster or the remains from an (already-disrupted) dwarf galaxy. Patchick 126 is already classified as in-situ GCs by Belokurov & Kravtsov (2023), suggesting that its formation occurred $\approx 11.7-12.7$ Gyr ago, according to in-situ GC ages. Therefore, we keep these lower limits, even if their ages seem to be older than those listed in Table 2. On the other hand, FS are more segregate at [Fe/H] $>-0.7$, while their ages span over a wide range of values including intermediate age with age $<6$ Gyr.
@@
Figure 7: Age-Metallicity distribution for FS (using catalogues by Bensby et al. (2017) as red points, and Casali et al. (2020) as cyan points) and for new GCs (orange points). On the top and right panels, we show the respective histograms and cumulative distributions for FS (red and cyan histograms and dashed black lines, respectively) and for new GCs (orange histograms and continuous black lines).
@@ 6 Discussion
The literature extensively addresses differences in metallicity distributions between FS and GC populations, not only within our Galaxy but also in external galaxies. For example, Forte et al. (1981) proposed that the colour variations observed in the spheroidal components of external galaxies (NGC 4374, NGC 4406, NGC 4621, M87) compared to their GC systems suggest a disparity in metallicity, indicating that GCs are more metal-poor than the spheroidal components. This implies a distinct chemical enrichment history for the GC system and the spheroidal component, suggesting that GCs likely form a component that is dynamically and chemically distinct from the spheroid. More recently, Larsen et al. (2012), in their study of the Fornax dwarf spheroidal galaxy’s GC system and its metal-poor FS, suggested that the primary distinction in the metallicity distribution between FS and GCs arises from the ability of each Fornax cluster to accrete a substantial amount of gas with the same composition as the first-generation stars. In contrast, limited gas of this composition was available for the formation of additional FS.
In this Section, we want to sum up and to be able to understand if there is a certain link between FS and GCs, and if this can help solve the puzzle and try to answer questions in Section 5.
@@
One caveat is that, our GC compilation may still be incomplete as many other star clusters could remain hidden, due to crowding and extinction. Furthermore, we predict that the chemical analysis of these clusters will be surely performed, in order to confirm/refute the nature of the new GCs included in the present work, as well as to better distinguish between GC and OC nature. Finally, we need further support from numerical simulations to understand whether it is possible to form super-metal-rich GCs, understand their probable evolution until their extinction, or on the other hand understand why it is not possible to form them.
@@
However, we find surprising that there are a high percentage of super metal-rich FS, while no one GCs in the same metallicity range. For instance, presuming that we are roughly counting that there are $30-40\%$ of stars with [Fe/H] ¿ 0, so out of $\sim 50$ bulge GCs we expect $\sim 20-25$ to have super-Solar metallicity, but there are none. This fact is independent of how one defines a bulge GC; because there is no definitive classification, the criteria used to describe them sometimes changes author-by-author. Again, if we use another sample of fewer bulge GCs, only 30 bulge GCs in total, but there is still zero super-Solar metallicity GC. Therefore, we foresee that the answer to all our questions must be sought in the evolutionary difference between these two stellar populations, since we found significant differences, as displayed in Figure 7, where we recognise three evolutionary moments. We can suppose that young (age $<5-6$ Gyr) and (super) metal-rich FS may be the result of a dissolution of old open clusters (OCs), since they are younger, more metal-rich, less massive, less bound, with shallower potential wells and therefore with shorter break-up times than GCs (with the caveat that the OC-GC boundary is not so well-defined). This may be supported by the fact that many OCs can disintegrate, as recently suggested by Bhattacharya et al. (2022). In fact, they found that many OCs show extended features, such as tidal tails and extended coronas.
@@
Similar features have also been discovered in MW GCs. For instance, Kundu et al. (2019, 2021) concluded that some GCs analysed in the MW halo show the presence of extra-tidal RR Lyrae stars. So, we can expect, and in the future explore, that several small and low-luminosity GCs here analysed may be show similar extra-tidal features, as found by Ferrone et al. (2023) in other Galactic GCs (e.g. NGC 3201, NGC 4590, NGC 5466, Pal 5). But if GCs’ extra-tidal features are composed by stars like RR Lyrae, these in particular contribute with old and metal-poor stars. Therefore, if we speculate on the fact that (super) metal-rich FS are the result of OC dissolution, or groups of stars formed later from [Fe/H]-enriched gas, probably we do not see super metal-rich GCs because at that time, when GCs form (10-13 Gyr ago), the gas did not have enough time to get rich in metals to form super metal-rich GCs, so we suggest that they never formed.
@@
On the other hand, Navarro et al. (2023) also model GC destruction for objects with orbits that bring them close to the Galactic centre, finding that the least massive clusters are quickly ($<10$ Myr) destroyed. Instead, the most massive clusters have a long evolution, showing variations in the morphology, especially after each passage close to the supermassive black hole. Therefore, in order to explain why we do not detect super metal-rich GCs, we cannot exclude other options: (i) we have not detected them yet, because they live in very crowded and reddened regions where the separation from FS is impossible, or (ii) they have been destroyed, because they lived in preferentially dangerous regions, or have orbits that are prone to efficient dynamical effects, or (iii) they did not form, because of some unknown reason, or (iv) another possibility is the accretion of metal-rich elliptical galaxy like M32, that is not known to contain GCs (Rudenko et al. 2009), or (v) finally, a combination of these tentative answers may be the correct way to explain why super metal-rich GCs were not observed today.

### v2

@@ 4 Metallicity distribution
Moreover, comparing all distributions, we can see that the number density of the metal-poor part is in better agreement, whereas the metal-rich one is shifted to higher values for the new and bulge GC samples than the other two. We can expect this effect because the metal-rich component is concentrated more toward the inner regions of the Galaxy, where the majority of our new targets are located, while the metal-poor one becomes more significant at larger Galactocentric distances. Indeed, as found by Tonini (2013) model (their Figure 2), the metal-rich distribution was formed locally in the main progenitor at $z\sim 2$, while the metal-poor component was formed in satellites at an epoch $z\sim 3-4$, and then merged with the main progenitor. Therefore, connecting to the discussion made in Section 3, and comparing our distribution with the Tonini’s model, the metallicity bimodaly is a direct prediction of a hierarchical clustering scenario. In consequence, we argue that the metal-poor component is mainly the contribution due to the tidal disruption of dwarf-like objects, whereas the metal-rich population is related to the formation of the bulge and/or disk (see e.g., Mouhcine 2006). This may imply that many of our targets may be formed in situ, and they might be classified as main progenitors, following the nomenclature from Massari et al. (2019).
@@
Moreover, the height of the MD’s peaks can give us clues about the star formation histories. Figures 5 and 6 show that metal-poor and metal-rich peaks are similar, especially if we include in the total compilation the new GC sample. Therefore, based on Tonini’s model, this trend suggests that the MW experienced a balanced star formation ($M_{SF}/M_{0}=0.2$) and merger history ($M_{1}/M_{0}=0.3$) [^6] Where $M_{0}$ is the final stellar mass, $M_{SF}/M_{0}$ star formation history parameter, and $M_{1}/M_{0}$ is assembly histories parameter.. Moreover, we know that multiple components, with different ages and metallicity, co-exist in the bulge, suggesting that star formation does not stop at $z\sim 2$; father, the final stellar mass is formed at later (or $z<2$) time, hinting that younger (e.g., age of $\sim 10$ Gyr) GCs may have formed later in the MW bulge, due to an increasing in gas and so a burst in star formation due to merging events. Once the age of our targets has been confirmed, our young candidates might belong to this younger populations, and this can help confirm whether the formation of GCs may occur later in the MW bulge, indicating a strong presence of gas at that stage; otherwise this could be linked to accretion events, which enriched our Galaxy and formed a new generation of star clusters. However, this is a speculation until this statement is confirmed observationally as well as via simulations and models.
@@
Finally, we want to highlight the fact that spectroscopic analysis should be performed to recover orbital parameters and chemical abundances, thus constraining the real nature of these systems. Furthermore, it is likely that many other GCs are missing from our compilation.
@@
Figure 6: Same details as in Figure 5 (right panel). We split the Harris (1996) (2010 edition) GC catalog into inner (bulge) GCs located at $R_{G}<3$ kpc (red line) and outer (halo) GCs at $R_{G}>3$ kpc (blue line). The new GCs sample is drawn in black line. We use this comparison to better understand the separation between bulge and halo clusters and to show that new GCs sample follow a similar trend as the inner GCs.
@@ 5 Comparisons with field stars
Another important comparison we carried out within the framework of this study considers GCs and FS MDs (see Figure 5). In order to do this, we took into consideration five bulge FS public samples by Zoccali et al. (2008); Bensby et al. (2017); Hasselquist et al. (2020); Abdurro’uf et al. (2022). For the Abdurro’uf et al. (2022) catalog, we applied the typical APOGEE selection cuts: apo_ps41 $<3.5$, ASPCAPFLAG $=0$, STARFLAG $=0$, SNR $>70$, LOGG $<3.6$, TEFF $<5500$, VSCATTER $<1$, and ruwe $<1.4$. Also, for comparison, we included Galactic disk FSs, using the catalog from Casali et al. (2020).
@@
Even though they come from widely different surveys (this was done intentionally since we need high statistics), these are all in good agreement, exhibiting metallicity peaks between -0.5 and 0.0, with a long tail toward the metal-poor regime. In addition, we can see that both bulge and disk FS catalogs seems to follow the same behavior. From Figure 5, we can see that $\sim 30-40\%$ of stars in all those catalogs show metallicities higher than Solar value, while a very low percentage of stars have [Fe/H] $<-1.0$. Even though a large number of very metal-rich stars have been observed, we still do not know of any GCs in the super-Solar metallicity regime today. So, we consider the question of where the super metal-rich bulge GCs reside and whether they can feasibly form and exist. Alternatively, we explore whether there might be a physical process that preferentially destroyed them in the past.
@@
To gain clarity in these areas, we need to understand if there is a connection between GCs and isolated FSs. We could resort to a range of possible explanations, such as different star formation scenarios that may have occurred for these two populations, in addition to scenarios related to the formation of the bulge and disk during the early phases of the Universe (Costantin et al. 2021). Evolutionary effects, such as cluster disruption, mass loss, interactions, merging events may also play a role (Minniti et al. 2017, 2019). It is likely that all of these properties and effects have had an effect on the ratio of clusters to FSs at a given metallicity range and location within the Galaxy, and may be responsible for these observed differences. To date, many studies were focused on the Galactic halo, finding a good agreement between observations and simulations. For instance, Reina-Campos et al. (2020), using an E-MOSAIC simulation, found that GCs play a sub-dominant role in the build-up of the stellar halo. Additionally, these authors found that the disruption of GCs contributes between $0.3-14\%$ of the bulge mass, in agreement with observational estimates by Hughes et al. (2019b). It was also predicted that stars from disrupted GCs in the Galactic bulge show a higher fraction around [Fe/H] $\approx-1$. A similar fraction was found by Schiavon et al. (2016) when analysing FSs in the MW bulge in a specific metallicity range of [Fe/H] $<-1$ and, by finding nitrogen-enriched stars, concluded that $14\%$ of the stellar mass of the bulge came from disrupted GCs. Therefore, the contribution of disrupted GCs to the stellar population of the bulge constitutes only a small percentage. Thus, we suspect that this factor alone cannot explain this inconsistency in metallicity between clusters and bulge FSs; it is more likely that the ratio of FSs to GCs varies strongly in different environments. For instance, as found by Lamers et al. (2017) the difference between the GC and FS metallicity distributions is not specific to our Galaxy, but is a well known phenomenon in external galaxies as well and it changes if outer or inner regions are taken into account. Reporting what they found: (i) in dwarf galaxies the population is generally metal-poor, but the GCs are usually more metal-poor than FSs; (ii) in the M31 inner and outer halo, the cluster-to-star ratio is high at low metallicity and low at high metallicity; (iii) in both the inner and outer regions the relative contribution of the clusters at low metallicity is larger than that of the stars; (iv) in both the inner and the outer regions, the relative contributions of the clusters at high metallicity is smaller than that of the stars; and (v) finally, the ratios of Fornax dwarf spheroidal galaxy are more similar to those in the outer regions of the larger galaxies than to the inner regions.
6 Age-metallicity distribution
@@
Another distribution that we want to take into consideration in order to understand the main differences between GCs and FSs is the age-metallicity distribution (Figure 7). We used the catalogs by Bensby et al. (2017) (red points) and Casali et al. (2020) (cyan points), while we considered only the new GCs (orange points) by the present compilation, since (in general), these are low-luminosity and low-mass clusters, so we foresee that they have lost a fraction of stars during their life (see Section 3). As expected, we can see that there is an anti-correlation between age and metallicity, since the age decreases with metallicity. In particular, we may note that the younger and metal-rich population is represented by FSs, while GCs occupy the right-bottom part of the diagram, with lower metallicities and older ages. This is also highlighted by the histograms and cumulative distributions displayed by Figure 7. GCs are more segregated at age $>10$ Gyr and their [Fe/H] spans over all values, especially at [Fe/H] $<-0.5$. In our compilation, the only two clusters with age $<10$ are VVVCL002 (age $>6.5$ Gyr) and Patchick 126 (age $>8$ Gyr). For the first, Minniti et al. (2023) demonstrated that VVVCL0002 is an old GC, since its high $\alpha$-enhancement ($[\alpha/Fe]>+0.4$) is connected to rapid chemical evolution dominated by core-collapse supernovae rather than type Ia supernovae, formed together with other clusters and field stars present today in the Galactic bulge, rather than a younger open cluster or the remnants from an (already-disrupted) dwarf galaxy. Patchick 126 is already classified as an in situ GC by Belokurov & Kravtsov (2023), suggesting that its formation occurred $\approx 11.7-12.7$ Gyr ago, according to in situ GC ages. Therefore, we keep these lower limits, even if their ages seem to be older than those listed in Table 2. On the other hand, FSs are more segregate at [Fe/H] $>-0.7$, while their ages span over a wide range of values including intermediate age with age $<6$ Gyr.
@@
Figure 7: Age-metallicity distribution for FSs (using catalogs from Bensby et al. (2017) as red points and Casali et al. (2020) as cyan points) and for new GCs (orange points). In the top and right panels, we show the respective histograms and cumulative distributions for FSs (red and cyan histograms and dashed black lines, respectively) and for new GCs (orange histograms and continuous black lines).
@@ 7 Discussion
The literature extensively addresses differences in metallicity distributions between FS and GC populations, not only within our Galaxy, but also in external galaxies. For example, Forte et al. (1981) proposed that the color variations observed in the spheroidal components of external galaxies (NGC 4374, NGC 4406, NGC 4621, and M87) compared to their GC systems suggest a disparity in metallicity, indicating that GCs are more metal-poor than the spheroidal components. This implies a distinct chemical enrichment history for the GC system and the spheroidal component, suggesting that GCs are likely form a component that is dynamically and chemically distinct from the spheroid. More recently, Larsen et al. (2012), in their study of the Fornax dwarf spheroidal galaxy’s GC system and its metal-poor FSs, suggested that the primary distinction in the metallicity distribution between FSs and GCs arises from the ability of each Fornax cluster to accrete a substantial amount of gas with the same composition as the first-generation stars. In contrast, limited gas of this composition was available for the formation of additional FSs. In this section, we want to sum up and attempt to understand whether there is indeed a verified link between FSs and GCs. We also explore how this might help answer the questions posed in Section 5.
@@
One caveat is related to the fact that our GC compilation may still be incomplete, as many other star clusters could remain hidden, due to crowding and extinction. Furthermore, we predict that the chemical analysis of these clusters will be performed in order to confirm or refute the nature of the new GCs included in the present work, as well as to better distinguish between the natures of GCs and OCs. Finally, we need further support from numerical simulations to understand whether it is possible to form super-metal-rich GCs, understand their probable evolution until their extinction, or (on the other hand) to understand why it is not possible to form them.
@@
However, we find it surprising that there are a high percentage of super metal-rich FSs, but no GCs in the same metallicity range. For instance, presuming that we are roughly counting that there are $30-40\%$ of stars with [Fe/H] ¿ 0, out of $\sim 50$ bulge GCs, we would expect $\sim 20-25$ to have super-Solar metallicity; however, there are none. This fact is independent of how the bulge GC is defined, namely: because there is no definitive classification, the criteria used to describe them can differ among works in the literature. Again, if we use another sample of fewer bulge GCs, with only 30 bulge GCs in total, there is still not a single super-Solar metallicity GC present. Therefore, we foresee that the answers to all our questions must be sought in the evolutionary difference between these two stellar populations. This is supported by the fact that we found significant differences, as displayed in Figure 7, where we recognise three evolutionary moments. We can suppose that young (age $<5-6$ Gyr) and (super) metal-rich FSs may be the result of a dissolution of old open clusters (OCs), since they are younger, more metal-rich, less massive, less bound, with shallower potential wells and therefore with shorter break-up times than GCs (with the caveat that the OC-GC boundary is not so well defined). This may be supported by the fact that many OCs can disintegrate, as recently suggested by Bhattacharya et al. (2022). In fact, they found that many OCs show extended features, such as tidal tails and extended coronas.
@@
Similar features have also been discovered in MW GCs. For instance, Kundu et al. (2019, 2021) concluded that some GCs analyzed in the MW halo show the presence of extra-tidal RR Lyrae stars. So, we can expect (and attempt to explore in the future) the notion that several small and low-luminosity GCs analyzed here may be exhibit similar extra-tidal features, as found by Ferrone et al. (2023) in other Galactic GCs (e.g., NGC 3201, NGC 4590, NGC 5466, Pal 5). However, if a GC’s extra-tidal features are composed of stars such as RR Lyrae, these, in particular, contribute with old and metal-poor stars. Therefore, if we speculate on the fact that (super) metal-rich FSs are the result of OC dissolution or groups of stars formed later from [Fe/H]-enriched gas, we probably do not see super metal-rich GCs because at the time when GCs formed (10-13 Gyr ago), the gas did not have enough time to get rich in metals to form super metal-rich GCs. Thus, we suggest that they never formed in the first place.
@@
On the other hand, Navarro et al. (2023) also modeled GC destruction for objects with orbits that bring them close to the Galactic centre, finding that the least massive clusters are quickly ($<10$ Myr) destroyed. Instead, the most massive clusters have a lengthy evolution, displaying variations in the morphology, especially after each passage close to the supermassive black hole. Therefore, in order to explain why we have not yet detected super metal-rich GCs, we cannot exclude other options: (i) we have not detected them yet and that may be because they live in very crowded and reddened regions where the separation from FSs is impossible or (ii) they have been destroyed, because they lived in preferentially dangerous regions, or they have orbits that are prone to efficient dynamical effects; alternatively, (iii) they did not form, because of some unknown reason, (iv) another possibility is the accretion of metal-rich elliptical galaxy, such as M32, which is not known to contain GCs (Rudenko et al. 2009); or (v) finally, a combination of these tentative answers may serve as an apt approach to explaining why super metal-rich GCs are not observed today.